How Do I Solve This Problem???
(n-5)^2=20

Answer by stanbon(48568)

Take the square root of both sides to get:
n-5 = 2sqrt5 or -2sqrt5
n= 5+2sqrt5 or n=5-sqrt5
Cheers,
Stan H.